mxddi3:

simplify the radical with imaginary numbers:

mxddi3:

\[\sqrt{-2} *\sqrt{-10}\]

mxddi3:

@Hero

Hero:

Hint: Rewrite the expression usnig \(i = \sqrt{-1}\)

Hero:

For example \(\sqrt{-7} = \sqrt{-1 \cdot 7} = \sqrt{-1}\sqrt{7} = i\sqrt{7}\)

mxddi3:

so then \[\sqrt{-2}= \sqrt{-1*2}=\sqrt{-1}*\sqrt{2}= i \sqrt{2}\] and the same thing for -10, right?

Hero:

Very good. Correct. Do the same for the other expression then simplify completely. It would be best to show all your work here.

mxddi3:

ok so for the -10 you do \[\sqrt{-10}= \sqrt{-1*10}= \sqrt{-1} *\sqrt{10}= i \sqrt{10}\] so then you have \[i \sqrt{2}*i \sqrt{10}\] and then that is \[i^2\sqrt{20 }\] is this right so far? because then i simplify the sqrt of 20 and i^2=-1

Hero:

Correct. Keep going.

mxddi3:

ok so then \[i^2\sqrt{20}\] the i^2 becomes -1 so we have \[-1\sqrt{20}\] and then \[\sqrt{20}= \sqrt{4}*\sqrt{5}= 2\sqrt{5}\] and then the 2 and the -1 are outside the radical so you combine them and then you get \[-2\sqrt{5}\] ??

Hero:

Correct.

mxddi3:

ty!